The area of a rectangle is (8x^2+8x+2). The width is (2x+1). What is the length of the rectangle?
Answer by mananth(16949)

area of a rectangle is (8x^2+8x+2).
The width is (2x+1)
..
(8x^2+8x+2)
8x^2+4x+4x+2
4x(2x+1)+2(2x+1)
(2x+1)(4x+2)
..
Area of rectangle = L * W
(2x+1)(4x+2)= L* (2x+1)
(2x+1)(4x+2)/(2x+1)= L
L= 4x+2
L=2(2x+1)